Sep 11, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:30 a.m. Eastern (until switch from DST and then return to 11:00am)

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Determine the effect of the mis-placing ch 0-8 of SiPM in downstream modules. (Elton)
  2. Add iron shields to top cosmic-ray counters (Elton)
  3. Investigate difference between north and south trigger rates (Andrei)

Tentative Agenda

  1. Announcements
  2. Time of WG meeting
  3. Collaboration Meeting
  4. Action Items
  5. FCAL HDFCAL log book
  6. BCAL HDBCAL log book
    1. Renewal of loan agreement
    2. GlueX-doc-2540 Shayne Gryba
    3. GlueX-doc-2539 Laura Teigrob
  7. Reconstruction and Simulation
  8. Calibration
  9. Any other business


Attendees: Manuel, Yi, Will, Elton (JLab); Matt (IU); Zisis, Teigan, Noemi (new graduate student), Andrei (UofR)

  1. Announcements
    • Yesterday there was a broad power outage on the accelerator site. Hall D was off for about 20 min. One compressor from the solenoid refrigerator was affected. One crate of bias voltages on the BCAL (downstream south) did not come back on and will need investigation.
  2. Time of WG meeting
    • Decision to start at 11:30 till DST, then revert to 11:00 am. If the meetings are running too long, we can consider starting again at 11:00.
  3. Collaboration Meeting
    • Report on FCAL and BCAL 30 min each. Suggest we start at 9:00 am. Zisis will communicate this to Curtis.
  4. Action Items
  5. FCAL HDFCAL log book
    • Manuel took data with the LED pulser to investigate the threshold behavior vs the sum of 3 consecutive channels.
    • Substantial feedback was given. Some of this discussion will be taken offline and presented at the L1 trigger meeting on Mondays (next is 9/22).
    • Cleanup of NIM electronics to meet fire codes on the FCAL platform is ongoing and will be converted to VME.
  6. BCAL HDBCAL log book
    1. TDC thresholds
      • Mark has taken data with LEDs to determine the TDC threshold vs pulse integral
      • Scans over LED bias voltage show that there is a broad range of pulse sizes for a fixed bias voltage.
      • The threshold was checked for 50, 100 and 250 mV.
      • The distribution of thresholds is quite broad and could be due to varying pulse shapes. Next will investigate the threshold behavior as a function of the pulse height.
      • Yi: Noted that the distribution of pulse integrals varies considerably from channel to channel. Mark: This could be due to cuttoffs (low and high) in the range of biases used in the data run.
    2. Renewal of loan agreement
      • Elton: will need to determine how to process new MOUs.
    3. GlueX-doc-2540 Shayne Gryba
      • The difference in trigger rates between north and south should be investigated and Andrei will work on this next week.
    4. GlueX-doc-2539 Laura Teigrob
  7. Reconstruction (Teigan)
    • Implementing code to obtain the timing from the flash ADCs
    • Obtaining certificates and finalizing installation of all packages. Checking that all code compiles and runs. Understanding output is next.
  8. Simulation
    • Andrei/Irina have updated the geometry of the cosmic ray simulation to include the correct position of the trigger counters as well as added the iron for the soleniod yoke
  9. Calibration
    • Noemi will start attending the calibration meetings and work on BCAL calibration
    • Will has started to use BGGEN to generate events and will investigate the amount of data needed to calibrate the BCAL with pi0s and/or etas.
    • This will help us develop a run plan.
    • Part of the reconstruction meeting next week will be devoted to this topic and Will can give an update.
  10. Any other business